OpenStudy (anonymous):

Please please help! An air freshener starts with 44 grams and evaporates. In each of the following cases, write a formula for the quantity, Q grams, of air freshener remaining t days after the start. (a) The decrease is 4 grams per day. Q(t) = (b) The decrease is 9% per day. Q(t) = (c) The decrease is at a continuous rate of 12% per day. Q(t) =

OpenStudy (lanhikari22):

well i'd guess, (a) Q(t) = 4(11-t) // since at day 1 it would be 4(10) = 40 , day 2 = 4(9) = 36 ... etc (b) Q(t) = 44( 91%^t ) // since Day 1 = 44 *91% , Day 2 = (44 * 91%) * 91% ... etc (c) Q(t) = 44 - (12%) * t // since i assumed that by "continuous" the problem meant that the decrease is constant to the original. I could be wrong though, I hope someone could double-check me or something. well that's what i thought of, I hope you can find your solution if this isn't it.

OpenStudy (aum):

(a) The decrease is 4 grams per day. In t days, 4 * t grams would have been lost. What is left is Q(t) = 44 - 4t = 4(11 - t) (b) The decrease is 9% per day. Use the "compound interest formula": Q(t) = 44( 1- 0.09)^t = 44(0.91)^t (c) The decrease is at a continuous rate of 12% per day. Every INSTANT it is decreasing not just at the end of each day. Use the "continuous compounding formula": Q(t) = 44e^(.12t)

OpenStudy (anonymous):

the first and second are right the third shows it's wrong

OpenStudy (lanhikari22):

well i figured my interpretation was pretty shaky, thus i asked for a double-check.

OpenStudy (aum):

c) \(\large Q(t) = 44e^{-0.12t}\)